express the area A of a right triangle as a function of the height h if the base of the triangle is twice its height?
Answer by mananth(16946)

height be h
base = 2h
Area of triangle = 1/2 * base * height
Area = 1/2*h*2h
Area = h^2
...
m.ananth@hotmail.ca